# Learning Monotonic Attention in Transducer for Streaming Generation

### v1

@@ 5.4 Results of Speech Generation
AL ($ms,\downarrow$) 153 -
Table 4: Performance of MonoAttn-Transducer with different choices of prior alignment.
Chunk Size ($ms$) 320 640 960 1280
$p^{\mathrm{dia}}$ BLEU ($\uparrow$) 24.72 26.74 27.05 27.41
AL ($ms,\downarrow$) 997 1239 1606 1991
$p^{\mathrm{uni}}$ BLEU ($\uparrow$) 24.89 26.68 27.26 27.11
AL ($ms,\downarrow$) 993 1249 1601 1983
@@ 6 Analysis
6.1 Choice of Prior Alignment
In Section 3.2.3, we introduced two choices for prior alignment: the uniform prior $p^{\mathrm{uni}}$, which assumes an equal probability of generation at each time step; and the diagonal prior $p^{\mathrm{dia}}$, which prefers ideal synchrony between the source and target. We employed the diagonal prior $p^{\mathrm{dia}}$ as the default choice in the aforementioned experiments. In this section, we examine the impact of different choices.The results are displayed in Table 4. As shown, MonoAttn-Transducer’s performance demonstrates robustness to the choice of prior alignment, with only minor impacts on both translation quality and latency across all chunk size settings. In Appendix D, we visualize the posterior alignment when using different priors. We have observed that, even with significant differences in the prior distribution, the posterior remains fairly robust when the chunk size is constant. This nice property reinforces the robustness of using the inferred posterior to train monotonic attention.
@@ 7 Conclusion
In this paper, we propose an efficient algorithm for Transducer models to learn monotonic attention. Extensive experiments demonstrate that our MonoAttn-Transducer significantly improves the ability in handling non-monotonic alignments in streaming generation, offering a robust solution for Transducer-based frameworks to tackle more complex streaming generation tasks.
@@ Appendix B Baselines
EDAtt (Papi et al., 2023): It calculates the attention scores towards the latest received frames of speech, serving as guidance for an offline-trained translation model during simultaneous inference.
@@ Appendix C Numerical Results
In addition to Average Lagging (AL; Ma et al., 2020c), we also incorporate Average Proportion (AP; Cho & Esipova, 2016), Differentiable Average Lagging (DAL; Arivazhagan et al., 2019) and Length Adaptive Average Lagging (LAAL; Papi et al., 2022) as metrics to evaluate the latency. AL, DAL and LAAL are all reported with milliseconds. The trade-off between latency and translation quality is attained by adjusting the chunk size $C$. The offline results are obtained by setting the chunk size to be longer than any utterance in the dataset ($C=\infty$). We use SimulEval v1.1.4 for evaluation in all the experiments. The numerical results of MonoAttn-Transducer are presented in Table 5 and 6. A comparison of the computation-aware latency metrics for AL and LAAL between the Transducer and MonoAttn-Transducer models is presented in Table 7.
@@ Appendix D Visualization
In this section, we present more examples of diagonal prior and its posterior from training corpus. Additionally, we also provide examples of uniform prior and its posterior for comparison. We have observed that, even with significant differences in the prior distribution, the posterior remains fairly robust when the chunk size is constant. The vertical axis represents the target subword sequence and the horizontal axis represents the speech waveform. Darker areas indicate higher alignment probabilities. We use Montreal Forced Alignment tools (McAuliffe et al., 2017) to obtain speech-transcription alignments for illustration.
@@
Figure 8: Chunk size in this example is set to 320ms. (Uniform Prior)
Figure 9: Chunk size in this example is set to 640ms. (Uniform Prior)
Figure 10: Chunk size in this example is set to 960ms. (Uniform Prior)
Figure 11: Chunk size in this example is set to 1280ms. (Uniform Prior)

### v2

@@ 5.4 Results of Speech Generation
LAAL ($ms$) 984 -
StartOffset ($ms$) 1520 -
MonoAttn-Transducer ASR-BLEU ($\uparrow$) 18.3 19.3
LAAL ($ms$) 918 -
StartOffset ($ms$) 1491 -
@@ 7 Conclusion
In this paper, we propose an efficient algorithm to handle the non-monotonicity problem in Transducer-based streaming generation. Extensive experiments demonstrate that our MonoAttn-Transducer significantly improves the ability in handling non-monotonic alignments in streaming generation, offering a robust solution for Transducer-based frameworks to tackle more complex streaming generation tasks.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix B Baselines
EDAtt (Papi et al., 2023a): It calculates the attention scores towards the latest received frames of speech, serving as guidance for an offline-trained translation model during simultaneous inference.
AlignAtt (Papi et al., 2023b): It exploits the attention information to generate source-target alignments that guide the model during inference. AlignAtt demonstrates superior performance for delays exceeding 2 seconds, whereas our study primarily focuses on scenarios involving delays shorter than 2 seconds.
@@
LLM-based Models
LSG (Guo et al., 2025): It enables the LLM to devise a streaming generation policy that balances latency and generation quality by referring a baseline policy.
@@ Appendix C Numerical Results
In addition to Average Lagging (AL; Ma et al., 2020c), we also incorporate Average Proportion (AP; Cho & Esipova, 2016), Differentiable Average Lagging (DAL; Arivazhagan et al., 2019) and Length-Adaptive Average Lagging (LAAL; Papi et al., 2022) as metrics to evaluate the latency. AL, DAL and LAAL are all reported with milliseconds. The trade-off between latency and translation quality is attained by adjusting the chunk size $C$. The offline results are obtained by setting the chunk size to be longer than any utterance in the dataset ($C=\infty$). We use SimulEval v1.1.4 for evaluation in all the experiments. The numerical results of MonoAttn-Transducer are presented in Table 4 and 5. A comparison of the computation-aware latency metrics for AL and LAAL between the Transducer and MonoAttn-Transducer models is presented in Table 6.
@@ Appendix D Visualization
In this section, we present more examples of diagonal prior and its posterior from training corpus. We have observed that, even with significant differences in the prior distribution, the posterior remains fairly robust when the chunk size is constant. The vertical axis represents the target subword sequence and the horizontal axis represents the speech waveform. Darker areas indicate higher alignment probabilities. We use Montreal Forced Alignment tools (McAuliffe et al., 2017) to obtain speech-transcription alignments for illustration.